Grade 10Algebra

If the product of zeros of the polynomial ax2-6x-6 is 4. find the value of a.

karthikeya
10 Years agoGrade 10

1 Answer

ronit
10 Years ago
ax2-6x-6=0
product of roots of a quadratic equation in the form ax2+bx+c=0 is c/a
(-6)/a=4
a=(-6)/4
a=(-3)/2